This is NA61/SHINE electronic logbook (read-only BACKUP!)

Run summary Shift logbook Run settings Beam logbook BPD logbook DAQ logbook DCS logbook Experts logbook Gas logbook TOF logbook Trigger logbook PSD logbook VD logbook Run coordination Detector Upgrade DRS logbook
  Logbook for trigger experts, Page 1 of 1  NA61/SHINE logo
ID Date Author Subject Text Attachments
  97   Sun, 12.07.2026 14:15:59 J.SzewinskiSettings and preparations for the 07/2026 run   printscreen_97_20260712-141712.pngprintscreen_97_20260712-142135.png 
  96   Tue, 17.03.2026 17:11:32 J.SzewinskiFinal trigger configuration for the hydrogen run 03/2026Tue, 17.03.2026 17:17:51 Final trigger configuration
for the hydrogen run 03/2026 
 Screenshot_2026-03-17_at_17-10-32_TDAQ-Control.pngprintscreen_96_20260711-185714.png 
  95   Mon, 9.03.2026 15:43:53 J.SzewinskiNew oscilloscope installed in the trigger rackMon,  9.03.2026 15:44:10 
New oscilloscope from the electronics pool
has been installed in the trigger rack:
 printscreen_95_20260312-164046.pngprintscreen_95_20260314-154004.pngprintscreen_95_20260314-154015.pngprintscreen_95_20260314-191240.png 
  94   Mon, 8.12.2025 09:17:33 J.SzewinskiTrigger shutdownTrigger crate has been turned off, S3 and
S1-4 signals has been switched back to their
original positions.
 20251208_091057.jpg20251208_091222.jpg 
  93   Sun, 7.12.2025 20:48:31 J.SzewinskiFinal trigger configuration used during PbPb 2025 runSun,  7.12.2025 20:50:35 JS: Final trigger
configuration used during PbPb 2025 run
 Screenshot_2025-12-07_at_20-49-42_TDAQ-Control.png 
  92   Tue, 2.12.2025 14:28:45 J.SzewinskiTrigger settings for PB run 12/2025 (MF on)Trigger settings for PB run 12/2025 (MF on)
 printscreen_92_20251202-143415.pngprintscreen_92_20251202-143458.pngprintscreen_92_20251202-143525.pngprintscreen_92_20251202-143555.pngprintscreen_92_20251202-143634.png 
  91   Sun, 30.11.2025 16:16:05 J.SzewinskiTrigger for Pilot data for Pb 2025   Screenshot_2025-11-30_at_16-17-15_TDAQ-Control.png 
  90   Sat, 29.11.2025 20:21:47 J.SzewinskiTrigger threshold tuningSun, 30.11.2025 09:37:14 S3 threshold 180
mV -> T2/T1 ~ 69%  (Attachment 4)
Sun, 30.11.2025 09:37:14 S3 threshold 150
 8x 
  89   Wed, 5.11.2025 12:12:31 Jaroslaw SzewinskiTrigger settings for proton run 11/2025Trigger configuration for the run:

T1 = S1 S1 #V0 #V1 CED6
 Screenshot_2025-11-05_at_12-19-09_TDAQ-Control.png20251105_123449_small.jpg20251105_123528_small.jpg20251105_123628_small.jpgprintscreen_89_20251105-134016.png 
  88   Sat, 5.07.2025 11:59:07 J.SzewinskiTrigger tunning for fragmentation 07/2025Sat,  5.07.2025 11:59:59 Changing S4 threshold
from 200 mv to 230 mv for fragmentation run
  
  87   Sat, 7.06.2025 14:54:21 Jaroslaw SzewinskiTrigger preparation for neutrino run 2025Name of trigger configuration: Neutrino_2025_06_S3p_V0
Due to bad discriminator channels, 2 channels
are swapped:
  
  86   Mon, 2.12.2024 11:47:49 Dominik RybkaNIM crate exchangeMon,  2.12.2024 11:48:24 NIM crate for Trigger
system in Counting House exchanged due to
possible failure in future.
  
  85   Thu, 12.07.2018 21:18:58 E. KapturBUSY and Trigger delaysThu, 12.07.2018 21:19:07 

We removed 6.5 ns from MT singal entry to
  
  84   Fri, 17.11.2017 17:31:03 E. KapturThreshold on S3 changed17.11.2017 17:30:17 EK: [COLOR=red]I have
decreased threshold on S3 from 1.64V to 1.55V.[/COLOR]
  
  83   Mon, 13.11.2017 19:38:13 E. Kaptur, A. Snoch19A GeV/c[B]S1:[/B]
Threshold V = -0.5127 V.
Threshold doesn't need to be changed.
  
  82   Mon, 6.11.2017 19:32:04 A. SnochTrigger setup for Xe+La 75A GeV/cThreshold S2 at maximum: 0.5166 V.
S2 voltage changed from 1136 V to 1030 V.
  
  81   Tue, 24.10.2017 03:55:24 A. Snoch, E. KapturTrigger setup1. Thresholds at S2, V0, and V1 were adjusted,
threshold at S3 not changed (due to the need
of data w/ and w/o the target).
  
  80   Tue, 24.10.2017 00:58:16 E. KapturChange of CFD channel for S11 signalDue to intermittent failure of the S11 CFD
channel I have changed S11 channel to S4
channel. S4 is disconnected. We need another
  
  79   Thu, 12.10.2017 14:40:12 A.Aduszkiewicztrigger timing tuningTrigger timings verified with help of Emil
S2 and S4 are OK
V0 is perfect, which means our recent intervention
  
  78   Tue, 10.10.2017 23:41:41 A.Marcinek,A.AduszkiewiczV0 +2nsWe found the V0 signal cable cut, probably
by the S2 support. The cable was 6ns, but
we could find only a 8ns one for replacement.
  
  77   Thu, 8.12.2016 13:03:00 Scaller readoutOutput from scaller
2016-12-08 12:49.37 INF: [main] running with
hardware access ($Revision: 5325 $)!
  
  76   Thu, 1.12.2016 21:00:42 E. Kaptur, P. PodlaskiS3 in visual scalerChanged channel 8 in visual scaler S4->S3   
  75   Thu, 17.11.2016 09:50:55 P. PodlaskiPSD threshold scanRemoved  delay from PSD signal before discriminator.
Rise time of signal from the PSD is long
- together with threshold timing changes.
  
  74   Tue, 15.11.2016 16:00:38 E. KapturTrigger cleaning51 cables not connected to anything were taken
out.
  
  73   Mon, 24.10.2016 09:29:46 Piotr PodlaskiSignals in patter units, scalers and ADCs pp400Pattern unit 1 (12 channels):
ch1 - MT
ch2 - T1prsc
 printscreen_73_20161106-050400.png 
  72   Sat, 22.10.2016 15:28:29 P. PodlaskiTrigger settings for pp@400 runV0 delay between splitter and CFD was changed
4ns->5ns
  
  71   Wed, 28.09.2016 21:56:03 E. KapturV1 delay changedV1 delay changed from 3ns to 2ns between LED
and delaybox
  
  70   Thu, 22.09.2016 00:56:37 E. KapturTrigger changes for 120 GeV/c protonsV1 was delayed by 2 ns between splitter and
LED
  
  69   Fri, 16.09.2016 13:31:55 E. KapturChange of camac delay box delay for CED8The signal delay (for pattern units and scalers)
for CED8 was changed to match signal delay
of (CED6 and CED7)
  
  68   Tue, 13.09.2016 16:07:08 E. KapturTrigger channel mappingTrigger:
Ch1: S1
Ch2: S2
  
  67   Fri, 9.09.2016 15:38:04 E. KapturPi+ beam at 60 GeV/cNo timing corrections needed.

10.09.2016 10:53:00 (EK) V1 signal is much
  
  66   Mon, 18.07.2016 05:32:54 Aleksandra Snoch, Emil KapturV0Cable V0-CFD was changed from 5 ns to 4 ns.   
  65   Fri, 15.07.2016 08:08:23 Aleksandra Snoch, Piotr Podlaski, Emil KapturTrigger configuration for p+Pb@80 GeV/c run[B]Changes:[/B]
-- V0 discriminator changed from the Leading
Edge Discriminator to the Constant Fraction
 trigg_pPb_80GeV.gif 
  64   Fri, 11.09.2015 14:37:32 sandrosignal naming / connectionsHere are the signals in the trigger

S1
  
  63   Thu, 10.09.2015 19:11:04 sandrofirst trigger for 400 GeV ppFirst trigger for pp@400 GeV setup with no
CEDAR (S5 is timed in)
all counters timed in
  
  60   Wed, 18.03.2015 22:36:11 Oskar Wyszynski30A GeV -> 40A GeV S1 ok
S2 ok
V1 -1ns
  
  59   Sat, 14.03.2015 20:19:08 Oskar WyszynskiV1 signal changedV1 -3ns   
  58   Tue, 10.03.2015 19:51:30 Oskar Wyszynski19A GeV -> 30A GeVS1    ok
S2    ok
V1   -1ns
  
  57   Tue, 3.03.2015 22:05:39 O. Wyszynski13A GeV -> 19A GeVS1 - ok
S2 - ok
V1 - ok
  
  56   Tue, 24.02.2015 16:24:00 O. WyszynskiMassive clean up27 cables were removed from: misc
3 cables were removed from: PSD
59 cables were removed from: Zdet and Adet
  
  55   Tue, 24.02.2015 15:39:27 O. WyszynskiSignals for Ar+Sc after clean-upPattern unit 1 (12 channels):
ch1 - MT
ch2 - T1prsc
  
  54   Sun, 15.02.2015 19:26:38 O. WyszynskiSignals for Ar+ScPattern unit 1 (12 channels):
ch1 - MT
ch2 - T1prsc
  
  53   Thu, 12.02.2015 17:33:53 O. WyszynskiPSD in ADCThe PSD signal was found in the ADC2, ch11.
The signal has been delayed by 40ns.
  
  52   Wed, 11.02.2015 23:15:47 O. WyszynskiChanges for Ar runS5 signal goes to C1/Z1 signal input 
and
PSD signal goes to Ced6/Z2 singal input of
 2015_ArCrossSection.cnf2015_ArProduction.cnfBaseConfig.cnf 
  51   Wed, 11.02.2015 22:07:35 A.LaszloInteraction rates in Ar+Sc (test)Target-in: https://na61-elog.cern.ch/Trigger+logbook/150211_220745/id51-2015-02-11-220528.png
Target-out: https://na61-elog.cern.ch/Trigger+logbook/150211_221505/id51-2015-02-11-221316.png
 id51-2015-02-11-220528.pngid51-2015-02-11-221316.pngid51-2015-02-11-221741.pngid51-2015-02-11-222409.png 
  50   Sun, 8.02.2015 21:38:47 O. Wyszynskina61trigThe ATI graphics card was removed (broken).
Now it is using the integrated one (intel).
  
  49   Sun, 14.12.2014 17:54:55 Z. FodorPSD signal timingWe checked the summed signal of PSD, it is
coming 80 ns after the S11 discriminated
signal on discriminator and
  
  48   Wed, 19.11.2014 22:01:50 O. WyszynskiTrigger cables delaysced6 - removed 80ns
ced8 - taken from fan-out
c1 - cleanup cabling as it is not use
  
  47   Mon, 17.11.2014 23:48:25 O. Wyszynski A. LaszloTrigger internal delay correctedDELAYb_2 was wrong.
Corrected from 12 to 18 cu.
It seems, that 2012 BeBe13 was wrong...
 2014pPbAt158GeVPerCVer2.cnf 
  46   Mon, 17.11.2014 18:14:58 O.Wyszynski, A.LaszloSignals connected to CAMAC beam/trigger acquisition in DAQ for 2014 158GeV/c p+Pb run (run18421-)Pattern unit 1 (12 channels):
ch1 - MT
ch2 - T1prsc
  
  45   Sat, 15.11.2014 18:06:00 O. WyszynskiCedar 7-foldCedar 7-fold signal to pattern unit delay
decreased by 80ns 
  
  44   Thu, 13.11.2014 22:58:56 A.LaszloTrigger config file for 2014 158GeV/c p+Pb
Located at: /home/online/TriggerCurrentConfigs/2014pPbAt158GeVPerC.cnf
on the na61trig.cern.ch machine
 2014pPbAt158GeVPerC.cnf 
  43   Tue, 11.11.2014 15:09:21 A. LaszloTrigger PMT HV settings for p+Pb runFor trigger:
U[S1_1] = -2200V
U[S2]   = -2080V
  
  42   Tue, 11.11.2014 14:31:47 O. WyszynskiThresholdsV0   294mV
V1   350mV ->  294mV
S4   120mV ->  104mV
  
  41   Tue, 11.11.2014 14:29:04 O. Wyszynski A. LaszloNew preliminary config for p+Pb   PreliminarySetup.cnf 
  40   Tue, 11.11.2014 13:42:16 O. WyszynskiTrigger internal delaysScan:
[TABLE border="1"]
D0|24|25|27
 BaseConfig.cnf 
  39   Sun, 2.11.2014 21:36:45 O. WyszynskiS4 signalRemove doubled attenuator from S4 signal.   
  38   Fri, 31.10.2014 19:54:40 O. WyszynskiThresholds and Voltages      new    (old)      new    (old)
S1_1  2200V            -330mV
S2    2000V  (1900V)   -100mV
  
  37   Wed, 23.01.2013 18:11:10 O. WyszynskiPattern Units[TABLE border="1"]
Channel| P.U. 1| P.U. 2
|-
  
  36   Tue, 15.01.2013 21:35:16 O.WyszynskiInternal delays changedScan:
[TABLE border="1"]
D0|22|23|24|25|27
  
  35   Mon, 10.12.2012 18:54:48 O.WyszynskiPSD in pattern unitSee: https://na61db.cern.ch:8443/elog/DAQ+logbook/37   
  34   Tue, 31.07.2012 21:42:28 O.WyszynskiModificationsPattern unit:
C1, CED6, CED7, CED8 was synchronized to
the gate.
  
  33   Tue, 31.07.2012 02:45:45 O.WyszynskiTrigger modificationsSignal S5 (delayed ~180ns) go to input B12
of Trigger.
Signal S1 go to input B13 of Trigger,
  
  32   Mon, 25.06.2012 06:08:42 A.LaszloSet beam only trigger configuration for PSD intensity testSet up beam only trigger with PID, 120GeV/c.

Config file: al_beamonlyPID_120.cnf
  
  31   Sun, 24.06.2012 12:52:44 Oskar WyszynskiConfig fileWe are using ow_beamPID_120.cnf.
T1 - s1, s2, v0, v1, ced6
T2 - T1 + s4bar
 id31-2012-06-24-131200.pngid31-2012-06-24-131210.png 
  30   Fri, 9.12.2011 13:18:30 H.-J. MathesLink to trigger software changedSwitched trigger software from Trigger.trunk.hmathes
(@5550) to v2011_11_19 (was created from
that trunk revision).
  
  29   Wed, 23.11.2011 20:47:21 S. PulawskiPb beam 80A GeV/cT4 scaled factor change to 20 for Pb beam.   
  28   Tue, 22.11.2011 02:55:37 S. PulawskiNew treshold for Be beam at 80A GeV/c Tresholds measure on discriminator:
Z2 -2.00V
Z3 -4.00 V
  
  27   Fri, 18.11.2011 14:00:31 sandrounprescaled triggersAdded a fan out for unprescaled triggers.
Please use this fan out
  
  26   Thu, 17.11.2011 13:12:15 sandromapping of pattern unitsPattern Unit 1
--------------
MT
  
  25   Thu, 17.11.2011 01:01:28 S. PulawskiTreshold for PSD set to 520change of treshhold.   
  24   Thu, 10.11.2011 10:09:37 H.-J. MathesTrigger setup S1*S2(HJM) File for trigger setup (S1*S2) is: config-13AGeV_S1S2.cnf

File for trigger setup (T1 = S1*S2*S4) is:
  
  23   Wed, 9.11.2011 11:29:42 H.-J. MathesTrigger electronics etc.10:00 (HJM) moved diagnostic scaler from Trigger
crate #3 (18-3) to Trigger crate #2 (18-2),
this should give space for PSD related trigger
  
  22   Wed, 9.11.2011 00:35:00 H.-J. MathesTrigger ...08.11 17:30 (HJM) uTrigger-Crate removed (=
cannibalized), now used for PSD and Zdet
trigger logic
  
  21   Fri, 4.11.2011 10:23:30 H.-J. MathesTrigger softwareSwitched to Trigger.trunk.hmathes, some ongoing
development and test...
  
  20   Fri, 16.09.2011 10:58:31 H.-J. MathesTrigger softwareHJM: Switched to Trigger.v2011_09_10, this
version (tag v2011_09_10) is identical to
the previously used one (trunk@4968),
  
  19   Fri, 16.09.2011 09:22:52 A.LaszloUsing S5 aircore cable for PSD trigger signalSee: https://na61db.cern.ch:8443/elog/Experts+logbook/91   
  18   Sun, 11.09.2011 11:53:40 H.-J.MathesTrigger softwareHJM 11:51: Switched trigger software from
v2011_08_31 to Trigger.trunk.hmathes (@4968).
  
  17   Wed, 31.08.2011 16:12:33 H.-J. MathesTrigger softwareSet symbolic link 'current' to v2011_08_31,
in case of problems (new ones), switch back
to v2011_08_02!
  
  16   Tue, 30.08.2011 16:06:55 H.-J. MathesTrigger software, switch to test version12:06 (HJM) running application readscalers2
stopped
            switched to test version in ~/Trigger/Trigger.trunk.hmathes
  
  15   Tue, 2.08.2011 19:11:38 H.-J. MathesSoftware link set to v2011_08_02Set symbolic link ~/Trigger/current to Trigger.v2011_08_02

This is trunk@4660 copied to tags@4661. Please
  
  14   Tue, 2.08.2011 19:03:46 H.-J. MathesSignal 'CLK' in the trigger system18:56 just discovered this: the signal 'CLK'
is no longer defined in signals.dat and also
not connected to ULM CMC206. Don't know when
  
  13   Tue, 2.08.2011 15:05:50 H.-J. Mathes, P. RoznoskiRerouting of V0p signal15:02: The routing of the signal from the
V0p counter was changed in order to shorten
the signal delay to the counting house:
  
  12   Tue, 2.08.2011 14:44:13 H.-J. MathesTrigger software update14:40: HJM - Updated to trunk@4658, this contains
Sandros changes (Monitor.cxx) and a (minor)
modification of the start/stop scripts.
  
  11   Sun, 31.07.2011 19:27:03 H.-J. MathesTrigger softwareUpdated to trunk@4634, the bug in handling
disconnected TCP/IP sockets is fixed now.
  
  10   Sun, 31.07.2011 05:19:05 sandrosetting up triggercounter   HV     thr.   comments
S1        2200   170    KFKI disc
S2        2000   50     KFKI disc
  
  9   Sat, 30.07.2011 10:49:01 roznowskibeam cnt settings change[CODE]
   | singal | treshold |
------------------------
  
  8   Fri, 29.07.2011 04:18:46 H.-J. MathesTrigger-Softwareon na61trig: Set link for current software
to Trigger.trunk (@4576)
  
  7   Wed, 27.07.2011 08:35:38 H.-J. MathesIcons on desktop of na61trig.cern.ch[H3]Trigger program icons for 2011 RUN[/H3]
HJM: For the 2011 RUN a new version of the
trigger software will be probably active
 id7-2011-07-27-111908.png 
  6   Fri, 13.08.2010 15:35:06 sandrotrigger checkchecked trigger setup for 13 GeV -> OK

checked trigger timing -> OK
  
  5   Sun, 8.08.2010 10:47:12 albertoTrigger configuration for Long target 158 GeV magnetic fieldSince S3 efficiency is very low at high field,
we changed trigger configuration as follows:
 id5-2010-08-08-102810.pngid5-2010-08-08-105917.png 
  4   Wed, 4.08.2010 18:00:26 sandros2 counting rateRUN 10850 between 17h00 and 17h30 for about
30 spills the scaler rates for S2 were bizarre.
Back to normal :-)
  
  3   Sat, 31.07.2010 18:21:54 Alain Blondelchange of trigger downscalingstudied trigger downscaling factors. 
T2 downscaling of 4:
s1 81835 
  
  2   Thu, 29.07.2010 18:02:39 alain blondel, sandro bravar, alberto marchionni trigger studiesTook a number of runs with specific trigger
configurations to check the alignment of
the counters, BPD, and target. 
 target-counters-alignment-29-07-2010.pdfbpd-10507.vi.pngBPDs10508.vi.pngBPDs10509.vi.pngBPDs10512.vi.png 
  1   Thu, 29.07.2010 03:32:05 Alain BlondelLong target triggerconsists of

T1 prescale 10   Cedar.(CKOV)bar.S1.S2.(V1')bar
  
ELOG V3.1.3-7933898